Question:


% Analyte = weight of Analyte/weight of sample × G.F ×100/1



G.F is the gravimetric factor = substance sought / substance known.


Calculations


1. Calculate the % sodium in a 0.721g detergent sample final Field is 0.320g NaP2O4.


2.A 0.2025 g sample consisting of only BaCl2 and KCl required 20.25 mL of 0.1200 M AgNO 3 solution for the quantitative precipitation of chloride.


Calculate the %K in the sample


1
Expert's answer
2021-02-09T03:45:57-0500

"\\% Analyte = \\dfrac{weight\\ of\\ Analyte}{weight of sample} \u00d7 G.F \u00d7\\dfrac{100}{1}\n\u200b"

"G.f=\dfrac{22.98}{148.93}"


"=0.15"

Substituting the value;

"\%Na=\dfrac{0.721}{0.320}*0.15*100\newline
\%of sodium =33.7"